### v1

# FinAudio: A Benchmark for Audio Large Language Models in Financial Applications

Abstract
Audio Large Language Models (AudioLLMs) have received widespread attention and have significantly improved performance on audio tasks such as conversation, audio understanding, and automatic speech recognition (ASR). Despite these advancements, there is an absence of a benchmark for assessing AudioLLMs in financial scenarios, where audio data, such as earnings conference calls and CEO speeches, are crucial resources for financial analysis and investment decisions. In this paper, we introduce FinAudio, the first benchmark designed to evaluate the capacity of AudioLLMs in the financial domain. We first define three tasks based on the unique characteristics of the financial domain: 1) ASR for short financial audio, 2) ASR for long financial audio, and 3) summarization of long financial audio. Then, we curate two short and two long audio datasets, respectively, and develop a novel dataset for financial audio summarization, comprising the FinAudio benchmark. Then, we evaluate seven prevalent AudioLLMs on FinAudio. Our evaluation reveals the limitations of existing AudioLLMs in the financial domain and offers insights for improving AudioLLMs. All datasets and codes will be released.
1 Introduction
Audio data plays a critical role in the financial domain, encompassing earnings conference calls, investor presentations, and customer service interactions (see Figure 1). Timely and accurate interpretation of financial audio data underpins financial services, sentiment analysis, and investment decisions Qin and Yang (2019); Yang et al. (2020); Cao et al. (2024).
Recent developments in financial AI have been propelled by breakthroughs in natural language processing (NLP), particularly large language models (LLMs), such as BloombergGPT Wu et al. (2023b), PIXIU Xie et al. (2023), FinGPT Liu et al. . Financial evaluation suites like FinBen Xie et al. (2025) further provide extensive assessments of LLMs in financial NLP tasks. Despite these advancements, there remains a critical gap: there is no audio-focused financial LLM or a corresponding evaluation benchmark. Although multimodal financial LLMs (e.g., FinLLaVA-8B Xie et al. (2024)) can process text, images, and charts, they cannot yet handle audio data.
Figure 1: An overview of critical financial audio data types and applications.
Figure 2: Overview of FinAudio benchmark.
In the general domain, AudioLLMs have made significant progress Wu et al. (2023a); Zhang et al. (2024). AudioLLMs extend foundational LLMs with an audio encoder that aligns audio inputs to the LLM’s input space Radford et al. (2023); Hu et al. (2024); Chu et al. (2023, 2024); Tang et al. , enabling audio-centric tasks such as automatic speech recognition (ASR), speech question answering (SQA), audio scene understanding, and voice analysis. While their effectiveness in these tasks has been explored in a variety of general scenarios Yang et al. (2024); Wang et al. (2024), a comprehensive benchmark for financial audio remains absent. This gap limits the research community’s ability to systematically evaluate performance, compare strategies, and identify biases in financial audio datasets, hindering downstream applications such as investment advisory and financial services.
To address this critical need, we introduce FinAudio, the first AudioLLM benchmark tailored for the financial domain. Specifically, FinAudio includes three tasks representating real-world financial scenarios: (1) ASR for short financial audio clips, (2) ASR for long financial audio recordings, and (3) summarization of financial audio data. For this purpose, we have collected four open-source financial audio datasets. Two of these datasets comprise numerous short audio clips (each under one minute), while the other two feature much longer recordings, about 45 to 60 minutes, which is sufficient to include a complete earnings conference call. Additionally, We collect a new dataset tailored for financial audio summarization, named FinAudioSum. Altogether, these five datasets form the FinAudio benchmark, comprising over 400 hours of financial audio. FinAudio’s scale is comparable to general-domain AudioLLM benchmarks.
On FinAudio, we evaluate seven representative AudioLLMs and observed the following: (a) AudioLLMs perform better in ASR tasks on short financial audio clips compared to longer recordings; (b) the performance of long-audio ASR limits audio summarization quality; (c) AudioLLMs exhibit significant variability in their robustness to different instructions; (d) The open-source AudioLLM Whisper-v3 outperforms closed-source models across all ASR datasets, offering a potential low-cost and privacy-preserving solution for both the financial industry and individual investors. Our main contributions are below:
(1) We propose FinAudio, the first comprehensive open-source evaluation benchmark for AudioLLMs in the financial audio data.
(2) We design three tasks that reflect realistic financial speech scenarios and compile five evaluation datasets, including a newly developed dataset for financial audio summarization.
(3) We conduct a systematic evaluation of seven AudioLLMs, showcasing their advantages and limitations and highlighting challenges and future research directions.
2 Related Work
2.1 Advancements in AudioLLM
Among the modalities addressed by multimodal LLMs, audio stands out as particularly critical and challenging Su et al. (2025). Several multimodal models, such as Gemini Pro Team et al. (2023) and PandaGPT Su et al. (2023), have been adapted to process audio data. Prominent AudioLLMs include AudioGPT Huang et al. (2024), SpeechGPT Zhang et al. (2023), AudioPaLM Rubenstein et al. (2023), Qwen-Audio Chu et al. (2023), WavLLM Hu et al. (2024), Whisper Radford et al. (2023), Salmonn Tang et al. , and Qwen2-Audio Chu et al. (2024). Correspondingly, general-domain benchmarks like AirBench Yang et al. (2024), emphasizing conversational scenarios, and AudioBench Wang et al. (2024), evaluating eight tasks across twenty datasets, have been developed to assess AudioLLMs. However, these benchmarks contain limited financial audio data, inadequate for a comprehensive evaluation in financial contexts.
Figure 3: Evaluation pipelines for the three FinAudio’s major tasks.
2.2 Progress in Multimodal Financial LLMs And Benchmark
Recent advances in general-domain LLMs have accelerated the emergence of specialized financial LLMs, such as FinGPT Liu et al. , InvestLM Yang et al. (2023), and BloombergGPT Wu et al. (2023b). The progression toward multimodal financial LLMs—including Open-FinLLMs Xie et al. (2024) and FinTral Bhatia et al. (2024)—further highlights efforts to address complex financial tasks involving charts and tabular data. These multimodal models overcome the limitations of traditional approaches Ang and Lim (2022); Bhatia et al. (2024); Yu and Suchow (2022) by more flexibly adapting to financial contexts. As financial LLMs become increasingly specialized, the demand for robust evaluation grows. In response, various benchmarks have been proposed. FLUE Shah et al. (2022) introduced the first benchmark for financial NLP evaluation, covering tasks such as sentiment analysis, headline classification, named entity recognition, structure boundary detection, and question answering. PIXIU Xie et al. (2023) extended these evaluations into multimodal document analysis, while FinBen Xie et al. (2025) significantly expanded coverage, featuring 36 datasets across 24 financial tasks. Despite these advances, existing benchmarks remain predominantly text-focused, overlooking audio—an essential modality in financial applications.
3 The FinAudio Benchmark
FinAudio aims to evaluate the performance of AudioLLMs on financial audio data. Our evaluation primarily focuses on the models’ ability to comprehend financial audio across various financial scenarios. Accordingly, our benchmarks are designed to evaluate three crucial aspects: 1) ASR for short financial audio clips, which are common in financial services and conversations; 2) ASR for long financial audio recordings, typically in major financial meetings; 3) Summarization for long financial audio recordings, aimed at testing the capacity of AudioLLMs to construct frameworks for understanding extensive financial audio content. For each task, we both compile existing datasets and develop new ones to ensure a comprehensive understanding ability. Figure 1 is an overview of FinAudio . This section will detail the task definitions and the development of the evaluation datasets.
3.1 FinAudio Tasks
Here we specify the FinAudio tasks, which are illustrated in Figure 3.
Dataset Name Type #Samples # Hours Task Metrics
MDRM-test Short Clips 22,208 87 short financial clip ASR WER
SPGISpeech-test Short Clips 39,341 130 short financial clip ASR WER
Earning-21 Long Audio 44 39 long financial audio ASR WER
Earning-22 Long Audio 125 120 long financial audio ASR WER
FinAudioSum Long Audio 64 55 long financial audio Summarization Rouge-L & BertScore
Table 1: Statistics of the datasets in the FinAudio benchmark.
3.1.1 ASR for short financial audio clip
Automatic Speech Recognition converts spoken language into text. Short audio clips from financial chatbot assistants or financial news often contain dense financial terms. This task evaluates AudioLLMs’ accuracy in transcribing financial audio clips, directly impacting applications such as financial voice assistants.
In this task, each input is represented as $(A,Q,R)$, where $A$ is an audio clip, $Q$ is the prompt instruction, and $R$ is the reference transcript. The transcribed text $T$ is generated as $T=\textit{AudioLLM}(A,Q)$. The Word Error Rate (WER) is computed as $\text{WER}=\frac{S+D+I}{N}$, where $S$, $D$, and $I$ represent the number of substitutions, deletions, and insertions, respectively, and $N$ is the total number of words in the reference ($N=S+D+C$, with $C$ indicating correct words). A lower WER indicates better ASR performance.
3.1.2 ASR for long financial audio recordings
Long audio recordings are common in finance, capturing events like quarterly earnings calls, mergers and acquisitions discussions, and central bank meetings. In this scenario, ASR task is thus used to transcribe lengthy financial audio, evaluating AudioLLMs’ capability to handle extended audio inputs. Challenges include long durations and complex financial terminology. This task assesses the models’ ability to consistently maintain transcription accuracy over long recordings, which is essential for comprehensive financial analysis.
To handle the varying window lengths and the inability of current AudioLLMs to process long audio files within a single window, we standardize audio inputs to 30-second segments, dividing each raw audio $A$ into chunks $\{a_{1},a_{2},\dots,a_{n}\}$. We transcribe each chunk $a_{i}$ using AudioLLMs into $t_{i}$, and concatenated the resulting transcriptions as $T=[t_{1};t_{2};\dots;t_{n}]$. The transcription quality is then evaluated by computing the WER.
3.1.3 Financial audio summarization
This task measures the ability of AudioLLMs to accurately summarize long financial audio content, a scenario unique to financial applications. It evaluates models’ understanding of financial discussions, extraction of key points, and effectiveness in condensing lengthy audio into concise summaries. This capability is crucial for stakeholders needing quick insights from extensive financial dialogues, such as investor briefings or regulatory meetings.
AudioLLMs can process audio with a maximum length of 30 seconds. Therefore, AudioLLM cannot perform the summarization task independently, we developed a processing pipeline. Given input data represented as $(A,Q,L)$, where $A$ denotes the long audio recording, $Q$ the prompt instruction, and $L$ the reference summary, the pipeline first segments the long audio into smaller chunks (see Section 3.1.2). Each chunk is transcribed by an AudioLLM into text segments $t_{i}$, which are combined into a complete transcription $T$. This transcription is then input into an LLM to generate the summary $S$. Summarization performance is evaluated using two widely adopted summarization evaluation metrics: Rouge-L Lin (2004) and BertScore Zhang et al. , computed between the generated summary ($S$) and the reference summary ($L$).
Datasets Models Whisper Qwen2-Audio Qwen2-Audio SALMONN SALMONN Gemini Gemini -v3 -7B -7B-Instruct -7B -13B -1.5-flash -2.0-flash MDRM-test 2.14 3.97 4.68 51.52 49.17 4.850 4.321 SPGISpeech-test 2.88 4.42 5.74 39.51 41.17 5.802 5.143 Earning-21 11.85 26.06 29.58 83.20 80.54 18.58 19.17 Earning-22 15.93 42.76 33.65 88.50 86.23 27.13 28.12
Table 2: Comparison of model performance based on WER across various speech datasets on ASR task.
3.2 FinAudio Datasets
The evaluation dataset in FinAudio is developed from two primary data sources: 1) existing open-source financial audio data originally created for non-LLM evaluation purposes. 2) novel datasets introduced in this work. Table 1 summarizes the key statistics. This section outlines their key information, leaving the details to Appendix A.
3.2.1 Short financial audio clip datasets
Two datasets provide financial audio clips: MDRM Qin and Yang (2019) and SPGISpeech O’Neill et al. (2021), both derived from earnings conference call recordings. The original audio data was segmented at the sentence level and aligned with corresponding transcripts. Each dataset is initially split into training and test subsets; we utilize only the test sets for evaluation. The MDRM-test set includes 22,208 audio clips totaling 87 hours, while the SPGISpeech-test set contains 39,341 clips totaling 130 hours.
3.2.2 Long financial audio recording datasets
Two datasets provide long financial audio recordings: Earnings-21 Del Rio et al. (2021), containing 44 earnings calls (approximately 39 hours), and Earnings-22 Del Rio et al. (2022), with 125 earnings calls (120 hours). Each recording includes a corresponding manual transcript. We use both datasets entirely for evaluation.
3.2.3 FinAudioSum
We create FinAudioSum based on the ECTSum dataset Mukherjee et al. (2022), originally designed for earnings call summarization using textual data. ECTSum comprises 2,425 earnings transcripts paired with expert-generated, telegram-style summaries. We obtain corresponding audio recordings for the ECTSum test set from earningcast [^1] https://earningscast.com/. Overlapping recordings with Earnings-21 and Earnings-22 (spanning 2019–2022) are removed. The final FinAudioSum dataset includes 64 recordings totaling 55 hours.
4 Model and Experiment Setup
We evaluate the performance of seven representative AudioLLMs on the FinAudio benchmark: five open-source models — Whisper Radford et al. (2023), Qwen2-Audio-7B, Qwen2-Audio-7B-instruct Chu et al. (2024), SALMONN-7B, SALMONN-13B Tang et al. ; two closed-source models — Gemini-1.5-flash and Gemini-2.0-flash Team et al. (2023). These models span various parameter sizes, open- and closed-source implementations, and include both base and instruction-tuned versions, enabling a comprehensive assessment of AudioLLMs for financial audio tasks. To ensure fairness, we standardized the prompt as: “Convert the audio speech into a text transcript,”. For summarization, transcripts from each model were summarized using GPT-4o (temperature=0). For open-sourced AudioLLMs, these experiments are exclusively conducted on two NVIDIA RTX A6000 GPUs, each equipped with 48GB of DRAM. Each round of evaluation takes approximately 200 hours to complete. For experiments in Tasks 1 and 2, we repeated evaluations three times and report the average results.
5 Results and Analysis
In this section, we first analyze the performance of seven models on two ASR tasks, then evaluate their summarization results, and finally examine their robustness to varying prompts.
5.1 Results on ASR Tasks
Results on short financial audio clips ASR.
Table 2 shows that Whisper-v3 achieves notably low WER scores (2%–3%) on both datasets, highlighting its robust speech recognition without domain-specific tuning. The Qwen2-Audio series (base and instruct versions) and Gemini exhibit moderate performance with WER scores between 4% and 6%. Conversely, SALMONN (7B and 13B) demonstrates substantially lower effectiveness, with WER around or exceeding 40%–50%. Despite SALMONN’s limited performance, these findings highlight the potential of other mainstream AudioLLMs to support voice-based financial services, particularly conversational assistants handling iterative short audio communications.
Figure 4: WER on models across different datasets. Whisper-v3 consistently achieved the lowest WER.
Results on long financial audio recording ASR.
In the long financial audio ASR task, Whisper-v3 remained the top performance. It achieves lower WER in the range of approximately 12% - 16%. The Gemini series emerged as the second-best performer, with WER spanning roughly 18–28%. The error rate of Qwen2-Audio series models has also increased significantly. SALMONN once again reported the highest error rates across both financial datasets (80–88%). These results reveal a decline in performance across all AudioLLMs, indicating that AudioLLMs face significant challenges when processing long financial audio data.
The SALMONN series models performed poorly on both ASR tasks. Upon reviewing the experimental logs, we observed that SALMONN frequently produced empty outputs or failed to understand prompt instructions. This suggests that SALMONN might be overfitting to specific audio data or tasks, lacks robust instruction-following capabilities, and exhibits a limited understanding of financial audio content.
In Figure 4, we find no significant difference in performance between Gemini-1.5 and Gemini-2.0. The open-source model Whisper-v3 achieves the best overall performance across all datasets. These results demonstrate robust capabilities even within long financial audio recognition. The Qwen-2 series, another open-source model, delivers comparable results to the closed-source Gemini models on short-audio ASR tasks but performs slightly worse on long-audio ASR. These findings demonstrate that open-source models can surpass closed-source counterparts in financial audio scenarios. This suggests that open-sourced AudioLLMs provide an effective, low-cost, and privacy-preserving solution for the financial industry and individual investors.
5.2 Results on summarization task
Dataset Whisper Qwen2-Audio-7B-Instruct Gemini-1.5-flash Rouge-L BertScore Rouge-L BertScore Rouge-L BertScore FinAudioSum 0.053 0.514 0.048 0.467 0.072 0.553
Table 3: Summarization performance using Rouge-L and BertScore metrics.
Given the high WER in the ASR tasks, effective summarization cannot be reliably performed by models with excessive transcription errors. Therefore, we excluded the SALMONN model from the summarization evaluation. Balancing model performance and computational cost, we selected three models—Whisper-v3, Qwen2-Audio-7B-Instruct, and Gemini-1.5-flash for subsequent summarization experiments.
As Table 3 shows, the summarization performance across all evaluated models is moderate. The overall summarization quality depends on the ASR performance of AudioLLMs. Gemini-1.5-flash achieves the highest scores in both ROUGE-L (0.072) and BERTScore (0.610), suggesting that Gemini’s transcription accuracy and semantic fidelity during the ASR stage align more closely with the ground-truth labels. Whisper outperforms Qwen2-Audio-7B-Instruct (ROUGE-L: 0.053 vs. 0.048; BERTScore: 0.514 vs. 0.467), demonstrating its robust general summarization capability even without fine-tuning on financial domain data.
5.3 Prompt Robustness Analysis
The previous experiments indicated that different models respond differently to the prompt. Therefore, we further conducted a prompt robustness analysis on three AudioLLMs: Whisper-v3, Qwen2-Audio-7B, and Qwen2-Audio-7B-Instruct - across three ASR datasets. Following the approach from AudioBench Wang et al. (2024), we constructed an instruction set comprising 10 prompts. These prompts convey similar instructions expressed in diverse formulations (see Appendix B). For each audio input, we randomly selected one prompt from this set to observe how variations in prompts affected model performance, as presented in Figure 5.
Figure 5: Prompt robustness analysis: comparison of WER between fixed-prompt and random-prompt trials.
We compare the WER in Figure 5 using two prompt types: the random prompt described above and the fixed prompt from Section 4, used in the main results (Table 2). Results for Qwen2-7B-Instruct and Whisper-v3 remain consistent, indicating their robustness across prompts. However, the performance of Qwen2-7B notably declines under varied prompts, exhibiting higher WER. The substantial performance gap between Qwen2-7B and Qwen2-7B-Instruct indicates that this AudioLLM is sensitive to prompt variations, providing empirical support for the necessity of instruction-tuning in AudioLLM training.
5.4 Error Analysis
During the ASR tasks, we noticed that AudioLLMs sometimes misinterpret numerical information and specialized financial terms in speech. Such errors significantly impact the factual accuracy of AudioLLMs within financial contexts. Specific examples of these errors are provided in the Appendix C.
6 Conclusion and Research Outlook
We propose FinAudio, an AudioLLM Benchmark for financial audio data. We constructed a total of 400h+ of audio data for three tasks and evaluated it on seven AudioLLMs. Future research directions for applying AudioLLMs in the financial industry include (1) extending the input context window length of AudioLLM, (2) enhancing the models’ comprehension of numerical transcription and specialized financial terminology, and (3) improving their reasoning capabilities in financial contexts.
Limitations
First, due to the limited availability of financial audio data, FinAudio primarily focuses on ASR-related tasks. Second, FinAudio does not address financial audio scenarios involving languages other than English.
Ethics Statement
All data sources used in FinAduio are public. The authors take full responsibility for the development of FinAduio, ensuring that the publicly available part of the dataset does not contain personal information and conforms to established ethical guidelines.
Appendix A Dataset Source
In section 3.2, we outline the statistical information for the datasets used in evaluation. Below, we provide a detailed description of the original datasets.
There are two datasets consisting of short financial audio clips:
  • MDRM Qin and Yang (2019): This is the first dataset to use multimodal financial data for stock volatility and movement prediction. MDRM includes both audio recordings and text transcripts from the 2017 earnings calls of 500 major companies listed on the S&P 500 and traded on U.S. stock exchanges. There are a total of 280 companies with 572 earnings conference calls. The author further segmented earnings conference calls into individual sentences based on the text and conducted corresponding audio segments. There are a total of 88,829 audio clips. We split the dataset into training and test sets using a 75%–25% ratio, with the test portion utilized as the evaluation set for the benchmark.
  • SPGISpecch O’Neill et al. (2021): The dataset comprises 5,000 hours of earnings conference call recordings. The original audio is segmented by sentence and professionally transcribed into a structured format. Each clip is between 5s and 15s long. The SPGISpeech dataset is originally split into training and test subsets; we utilize only the test subset for evaluation.
The combined test sets from the two datasets described above contain a total of 61,549 clips. This scale is sufficient for the evaluation of model performance. The corresponding training sets can be utilized for further supervised fine-tuning of AudioLLMs.
Additionally, there are two other publicly available earnings conference call datasets containing raw audio recordings, which can be leveraged for the long financial audio ASR task.
  • Earnings-21 Del Rio et al. (2021): The dataset comprises 44 complete earnings conference call recordings from 2021. A distinguishing characteristic of Earnings-21 is its emphasis on named entities, financial jargon, and numerical data, which often pose difficulties for ASR models. The aligned transcripts are further enhanced with detailed formatting and entity tags, making it possible to conduct granular error analysis.
  • Earnings-22 Del Rio et al. (2022): This is a updated version of Earnings-21. It has 125 audio samples with a total of 120 hours.
Both datasets were utilized for evaluating model performance.
Appendix B Instruction Set
As described in Section 5.3, we established an instruction set comprising 10 distinct prompts to evaluate the robustness of AudioLLMs. These prompts convey the same meaning. The detailed prompts are listed as follows:
  • (1) Convert the audio speech into a text transcript.
  • (2) Transcribe the spoken words into written form.
  • (3) Listen to the audio and provide the text version.
  • (4) Transform the speech into a text document.
  • (5) Capture the spoken language and convert it to text.
  • (6) Decode the audio and give me the written transcription.
  • (7) Recognize the verbal communication and transcribe it into text.
  • (8) Turn the speech input into a text transcription.
  • (9) Process the audio speech and provide the text output.
  • (10) Translate the spoken conversation into written text.
To assess the robustness of AudioLLMs to variations in input instructions, we randomly select a instruction from the above instruction set for each audio input. This approach closely simulates real-world scenarios, where user commands often differ in wording yet convey the same intent.
Appendix C Error Analysis
The experimental results demonstrate that ASR performance significantly influences downstream tasks, such as summarization. Through detailed analysis of the outputs, we identified the following major error categories:
  • Financial terminology error: Financial audio frequently contains specialized financial proper nouns, leading AudioLLMs to produce translation errors. For example, the company name “NextEra Energy" was incorrectly transcribed as “Era Energy."
  • Numerical information error : Financial audio typically includes extensive numerical information, such as monetary amounts. During transcription, AudioLLMs frequently encounter issues like digit inconsistencies or missing monetary units.